# pnasresearcharticle

### v1

@@ Abstract
This manuscript was compiled on February 7, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
5]3
@@
$Q=\frac{1}{2m}\sum_{k=1}^{K}\sum_{i,j\in C_{k}}\left[A_{ij}-\mathbf{E}(A_{ij}|\mathcal{H}_{0})\right],$ (1)
where $m$ is the total number of edges, the elements of adjacency matrix $A_{ij}$ contain the edge weights between nodes $i$ and $j$, and $\mathbf{E}(A_{ij}|\mathcal{H}_{0})$ expresses the expected proportion under the null hypothesis. The configuration model is a common null hypothesis that redistributes the edges weights evenly over the nodes; i.e., $k_{i}k_{j}/(2m)$ where $k_{i}$ is the degree of node $i$. Modularity can thus be seen as a statistical measure of the unexpectedness in edges arrangement and of how graph partitions are exploiting such modular structures to form densely connected (assortative) or bipartite (dissortative) communities (13, 14, 15). Many approaches have been proposed to optimize modularity, such as the Louvain algorithm which operates in the graph domain (16, 17) or the spectral method that eigendecomposes the modularity matrix ${\mathbf{B}}$ where $B_{ij}=A_{ij}-{k_{i}k_{j}}/(2m)$ (18, 19).
At first sight, the definition of modularity is amendable for directed graphs. First, the adjacency matrix considered in $Q$ can be asymmetric. Second, the null model can be changed to $k^{\text{out}}_{i}k^{\text{in}}_{j}/m$ to account for out- and in-degree of the nodes (20, 21). However, this type of directed modularity does not consider edge direction (22, 23); i.e., the contribution to modularity of a specific edge between nodes $i$ and $j$ remains constant as long as $A_{ij}+A_{ji}$ and the in- and out-degrees are the same (23). In addition, the spectral method cannot be applied to non-symmetric matrices.
Considerable work has been carried out to go beyond these limitations for community detection in directed networks (21, 23, 9). First, the spectral method has been applied to the symmetrized matrix $\mathbf{B}+\mathbf{B}^{T}$. However, because this approach obfuscates edge direction (21), the directed modularity has instead been redefined with random walk priors in the LinkRank method (23). It is defined as the difference between the fraction of time a random walker will spend within communities and the expected value of this fraction. While this approach accounts for edge direction and is coherent with the undirected modularity, the matrix definition of the LinkRank modularity is, however, asymmetric and the authors therefore present an optimization process that again requires symmetrization similar to the one in (21). Alternatively, the concept of co-clustering has been proposed (9). Essentially, the regularized graph Laplacian ${\mathbf{L}}$ of the directed graph (24) is computed and spectral clustering is then applied to the leading eigenvectors of $\mathbf{L}\mathbf{L}^{T}$ and $\mathbf{L}^{T}\mathbf{L}$ leading to two sets of clusters driven by outgoing and incoming edges, respectively. This approach does not integrate the notion of a null model proper to community structure and there is no natural correspondence between the clusters. In sum, there is not yet a satisfactory extension of community detection to directed graphs.
@@ Results
$Q_{\text{bi}}$ $=\frac{1}{m}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i\in C_{k}^{\text{out}}\\ j\in C_{k}^{\text{in}}\end{subarray}}\left[A_{ij}-\mathbf{E}(A_{ij}|{\mathcal{H}}_{0})\right],$ (2)
@@
Convex relaxation of the vectors $\mathbf{s}^{\text{out}}$ and $\mathbf{s}^{\text{in}}$ does not restrict their values to +1/1+1/-1+ 1 / - 1 and instead imposes a unit norm, which allows to obtain the solution by the singular value decomposition (SVD) $\mathbf{B}=\mathbf{U\Sigma V}^{T}$; see Methods for the full derivation. The columns of $\mathbf{U}$ contain the left singular vectors, those of $\mathbf{V}$ the right singular vectors, and the diagonal elements of $\mathbf{\Sigma}$ the corresponding singular values that relate to bimodularity up to a factor $m$.
Akin to Laplacian spectral embedding, different pairs of singular vectors $(\mathbf{u}_{k},\mathbf{v}_{k})$ provide representations of the network by decreasing bimodularity (25, 1). For the running example, the singular values are plotted in Fig. 1.C and reveal two components with large bimodularity. The visualization of the first component in Fig. 1.D essentially unveils the diagram of the underlying model, with the four sets clearly separated and edges running counter-clockwise.
Figure 2: Detection of bicommunities. A. $N$ left $\mathbf{u}$ and right $\mathbf{v}$ singular vectors of the SVD of the modularity matrix $\mathbf{B}$ (bimodularity embeddings). B. Aggregation of embeddings at the level of graph nodes to build a feature vector that represents the sending and receiving partition of the source ($i$) and target ($j$) node respectively. C. Feature matrix with features of all graph edges. D. Edge feature space (first two SVD components). Each dot represents an edge of the graph colored based on its corresponding cluster. E. Graph adjacency matrix where each edge cluster is colored. Sending (left) and receiving (top) nodes are shown with colors of the corresponding cluster. F. Bicommunities of the block-cycle graph with colors corresponding to the edge cluster. Colors and markers of nodes indicate whether a node belongs to the sending (red, square), receiving (blue, diamond) or both (circle, purple) community.
@@
Bimodularity-based embedding of the graph nodes provides insightful representations of network organization. However, an additional step is needed to effectively detect the bicommunities in terms of pairs of sending $C_{k}^{\text{out}}$ and receiving communities $C_{k}^{\text{in}}$.
@@
Figure 3: Singular value decomposition (SVD) of the modularity matrix built from the C. elegans wiring network of chemical synapses. (A) Spectrum of the singular values of the SVD. Colors indicate the sign of the singular value associated to assortative (red, $\mu>0$) or dissortative (blue, $\mu<0$) community structures. (B) Projection of the graph nodes (neurons) onto the left ($\mathbf{u}_{1}$, horizontal axis) and right ($\mathbf{v}_{1}$, vertical axis) singular vectors of the bimodular component associated with the largest singular value ($\mu_{1}=10.98$). Colors and shape of nodes indicate the type of the neurons (sensory: red square; interneuron: blue circle; or motor: green diamond). Edge colors indicate the type of the target node with the same color code as for the neuron types. Neuron labels are shown for graph nodes that are projected far from the origin $(0,0)$ of the axes.
@@
Figure 4 shows the 3 bicommunities of the C. elegans with highest bimodularity index (out of 5) extracted from the 5 first components of the SVD of $\mathbf{B}$. These clusters highlight specific information pathways in the worm’s motion system. In detail, $C_{3}$ identifies communication from body sensory processes to key interneurons in the worm’s brain (AVA, AVB, AVD). $C_{4}$ describes feed-forward transmissions of information from these interneurons to motor neurons of the body and interneurons of the tail (PVC). $C_{2}$ finally shows interconnectivity between motor neurons of the body and the neck as well as their communication with interneurons of the worm’s tail. Finer definitions of the C. elegans’s bicommunities is achieved with a larger number of edge clusters, see SI Appendix, Fig. S5.
Figure 4: Bicommunity of the C. elegans extracted from the $k=5$ first bimodularity embeddings (components of the SVD). Only the clusters with top 3 highest bimodularity index (out of 5) are displayed. A. The sending and receiving parts of each bicommunity are highlighted on the spatial location of neurons. Markers indicate whether a neuron belongs to the sending (red square), the receiving (blue diamond) or both (pink circle) parts of a bicommunity. The colors represent the difference between the proportion of edges in the sending and in the receiving partitions with red and blue for nodes that tend to send and receive more respectively. B. The distribution of neuron types (sensory, inter, motor and sex-specific neurons) in each sending or receiving community is summarized in the pie charts (middle). C. The spatial distribution of bicommunities along the tail to head axis is detailed in the histograms (right) as a proportion of the total number of neurons (gray bars). The sending and receiving patterns are separated in the upper and lower bars respectively. Colors indicate the type of neuron following the color scheme of the pie charts. The body and tail histograms are made bigger (twice their original size/number of neurons) for visualization purpose.
@@ Discussion
The application of the proposed scheme to the C. elegans’s neuronal wiring diagram further confirmed its relevance. In particular, the embedding according to the first component of the SVD encompasses the organization of the worm’s head and body locomotion system and untangles information pathways such as head-to-head, body-to-body and head-to-body sensory-motor flows. Specifically, this latter communication scheme is mainly driven by the AVE neurons that could integrate information from the sensory/control neurons of the head system and further transmit signals to the body control and motor neurons. We also observed that most body motor neurons also send information to more anterior motor neurons in the neck and the head, which confirms previous findings of intricate feedforward loop between the head and body motion systems (27, 28, 29). In short, the innovative aspect of bimodularity is that key clusters of neurons and their projections are captured in a single bimodular component (i.e., by a left/sending and right/receiving singular vectors) of the directed modularity matrix.
@@
There is one methodological aspect that needs to be addressed when dealing with SVD; i.e., the sign ambiguity singular values and vectors (30). Indeed, in the $\mathbf{B=U\Sigma V}^{T}$ decomposition the sign of a singular value and one singular vector: $\mathbf{B=U\Sigma V}^{T}=\mathbf{U(-\Sigma)(-V}^{T})$ can be freely swapped , which raises the need for a clear sign convention for univocal interpretation. In the case of undirected graph, the modularity matrix is symmetric and the singular value (eigenvalue) hints whether a community structure is assortative (positive) or rather dissortative (negative, also known as anti-modular) (14, 15, 31). Here, we propose a sign convention associated to the dot product between the two singular vectors (i.e., cosine of the angle between them) and show that it is related to the assortative (or dissortative) nature of the partitions, see SI Appendix, Fig. S7. In particular, we observe that the bimodularity index of a bicommunity is negative when there are more edges between the sending and receiving sets of nodes than within them. Therefore, this sign convention provides additional information about the observed community structure. In the C. elegans wiring network, the second component with the largest singular value is dissortative and supports the findings of the first component (Fig.3) by highlighting head interneurons that densely send information to body motor neurons without necessarily being connected (see SI Appendix, Fig. S8 for detailed description).
@@
While bimodularity is a suitable metric for highlighting community structure in directed networks, as an extension of modularity it shares some of its limitations such as the resolution limit (32). Future work should investigate how this limit is influenced by directed information. In Supplementary Material (SI Appendix, Fig. S9), we present a number of variations on the generated graphs that confirm the relevance and robustness of the bimodularity and bicommunity detection. The presented method for bicommunity detection is parametrized by the number of bimodular components $N$ and the number of desired clusters $K$. While $K$ is here tied to the clustering algorithm (i.e., k-means), we show that influence of $N$ is attenuated by multiplying entries in the edge feature (singular vectors $\mathbf{u}_{k}$ or $\mathbf{v}_{k}$) by their corresponding singular value $\mu_{k}$ (see SI Appendix, Fig. S2). This is a common approach in spectral clustering to incorporate the spectrum of eigen-/singular values to weight the matrix of eigen-/singular vectors that is then fed to clustering (25, 26).
@@
All code and data used in this article are openly available at https://github.com/MIPLabCH/Bimodularity.
@@
A directed binary graph was built from the wiring diagram of the C. elegans’s nervous system using information about 2194 chemical synapses between 279 neurons (graph nodes) (1). Out of all these connections, 1961 (89%) were asymmetric while only 233 (11%) were bidirectional. The undirected electric gap junction network was not used in the main analyses, but the bimodularity optimization and singular vector embedding of the joint (chemical synapses and gap junctions) graph are shown in SI Appendix, Fig. S10. Details about the type of each neuron (sensory, inter- or motor neurons) and their position in the worm’s body have been gathered from (29, 33) and are summarized in SI Dataset S1.
@@
$Q_{\text{bi}}$ $=\frac{1}{m}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i\in C_{k}^{\text{out}}\\ j\in C_{k}^{\text{in}}\end{subarray}}\left[A_{ij}-\mathbf{E}(A_{ij}|\mathcal{H}_{0})\right]$
@@
We consider the graph partition problem to separate the graph into two communities of nodes (sending community) that have common targets (receiving community). Let $\mathbf{s}^{\text{out}}$ be a separator vector that takes value $s_{i}^{\text{out}}=1$ if node $i$ is in one sending partition and $s_{i}^{\text{out}}=-1$ if it is in the other. Similarly, we define $\mathbf{s}^{\text{in}}$ as a separator vector for the receiving partition corresponding to the sending one. Therefore, $\delta_{C^{\text{out}},C^{\text{in}}}(i,j)$ can be rewritten as $\frac{s_{i}^{\text{out}}s_{j}^{\text{in}}+1}{2}$ that is $1$ if $s_{i}^{\text{out}}=s_{j}^{\text{in}}$ and 00 otherwise. We then develop the expression of bimodularity as:
@@
$=\frac{1}{2m}\sum\limits_{i,j}\left[A_{ij}-\frac{k_{i}^{\text{out}}k_{j}^{\text{in}}}{m}\right]s_{i}^{\text{out}}s_{j}^{\text{in}}+\frac{1}{2m}\sum\limits_{i,j}\left[A_{ij}-\frac{k_{i}^{\text{out}}k_{j}^{\text{in}}}{m}\right].$
Knowing that $\sum\limits_{i,j}A_{ij}=m$ ,that $\sum\limits_{i}k_{i}^{\text{out}}=m$ and that $\sum\limits_{j}k_{j}^{\text{in}}=m$, the second sum of the equation is equal to zero and the bimodularity index is can be expressed as a matrix multiplication:
@@
Maximization of bimodularity is achieved by looking for separator vectors $\mathbf{s}^{\text{out}},\mathbf{s}^{\text{in}}$ that maximize the bimodularity index $Q_{\text{bi}}(\mathbf{s}^{\text{out}},\mathbf{s}^{\text{in}})$ under the constraints that $\mathbf{s}^{\text{out}}$ and $\mathbf{s}^{\text{in}}$ are unitary vectors $(\mathbf{s}^{\text{out}})^{T}\mathbf{s}^{\text{out}}=1$ and $(\mathbf{s}^{\text{in}})^{T}\mathbf{s}^{\text{in}}=1$. We then write the unconstrainted optimization problem using Lagrange multipliers $\mu^{\text{out}}$ and $\mu^{\text{in}}$:
@@
Multiplying the first expression by $\mathbf{B}^{T}$, we then obtain $\mu^{\text{out}}\mathbf{B}^{T}\mathbf{s}^{\text{out}}=\mathbf{B}^{T}\mathbf{Bs}^{\text{in}}$, where we can substitute the second expression to obtain $\mu^{\text{out}}\mu^{\text{in}}\mathbf{s}^{\text{in}}=\mathbf{B}^{T}\mathbf{Bs}^{\text{in}}$. Here we recognize that $\mathbf{s}^{\text{in}}$ is an eigenvector of the symmetric matrix $\mathbf{B}^{T}\mathbf{B}$ with the corresponding eigenvalue $\lambda^{\text{in}}=\mu^{\text{out}}\mu^{\text{in}}$. Similarly, applying the same approach to the second (lower) expression and substituting the upper one shows that $\mathbf{s}^{\text{out}}$ is an eigenvector of $\mathbf{B}\mathbf{B}^{T}$ with the eigenvalue $\lambda^{\text{out}}=\mu^{\text{out}}\mu^{\text{in}}$.

### v2

@@ Abstract
This manuscript was compiled on July 21, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
3]3
@@
$Q=\frac{1}{2m}\sum_{k=1}^{K}\sum_{i,j\in C_{k}}\left[A_{ij}-\mathbb{E}(A_{ij}|\mathcal{H}_{0})\right],$ (1)
where $m$ is the total number of edges, the elements of adjacency matrix $A_{ij}$ contain the edge weights between nodes $i$ and $j$, and $\mathbb{E}(A_{ij}|\mathcal{H}_{0})$ expresses the expected proportion under the null hypothesis. The configuration model is a common null hypothesis that redistributes the edges weights evenly over the nodes; i.e., $k_{i}k_{j}/(2m)$ where $k_{i}$ is the degree of node $i$. Modularity can thus be seen as a statistical measure of the unexpectedness in edges arrangement and of how graph partitions are exploiting such modular structures to form densely connected (assortative) or bipartite (dissortative) communities (13, 14, 15). Many approaches have been proposed to optimize modularity, such as the Louvain algorithm which operates in the graph domain (16, 17) or the spectral method that eigendecomposes the modularity matrix ${\mathbf{B}}$ where $B_{ij}=A_{ij}-{k_{i}k_{j}}/(2m)$ (18, 19). Such optimizations can be interpreted as embedding problems for the graph nodes or edge, which in the latter may provide overlapping groups of nodes (20, 21).
At first sight, the definition of modularity is amendable for directed graphs. First, the adjacency matrix considered in $Q$ can be asymmetric. Second, the null model can be changed to $k^{\text{out}}_{i}k^{\text{in}}_{j}/m$ to account for out- and in-degree of the nodes (22, 23). However, this type of directed modularity does not consider edge direction (24, 25); i.e., the contribution to modularity of a specific edge between nodes $i$ and $j$ remains constant as long as $A_{ij}+A_{ji}$ and the in- and out-degrees are the same (25). In addition, the spectral method cannot be applied to non-symmetric matrices.
[…]
@@ Results
$Q_{\text{bi}}$ $=\frac{1}{m}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i\in C_{k}^{\text{out}}\\ j\in C_{k}^{\text{in}}\end{subarray}}\left[A_{ij}-\mathbb{E}(A_{ij}|{\mathcal{H}}_{0})\right],$ (2)
@@
Convex relaxation of the vectors $\mathbf{s}^{\text{out}}$ and $\mathbf{s}^{\text{in}}$ does not restrict their values to +1/1+1/-1+ 1 / - 1 and instead imposes a unit norm, which allows to obtain the solution by the SVD: $\mathbf{B}=\mathbf{U\Sigma V}^{T}$; see Methods for the full derivation. The columns of $\mathbf{U}$ contain the left singular vectors, those of $\mathbf{V}$ the right singular vectors, and the diagonal elements of $\mathbf{\Sigma}$ are the corresponding singular values that relate to bimodularity up to a factor $m$.
Akin to Laplacian spectral embedding, different pairs of singular vectors $(\mathbf{u}_{k},\mathbf{v}_{k})$ provide representations of the network by decreasing bimodularity (30, 1). For the running example, the singular values are plotted in Fig. 1.C and reveal two components with large bimodularity. The visualization of the first component in Fig. 1.D essentially unveils the diagram of the underlying model, with the four sets clearly separated and edges running counter-clockwise.
Figure 2: Detection of bicommunities. (A) $N$ left $\mathbf{u}$ and right $\mathbf{v}$ singular vectors of the SVD of the modularity matrix $\mathbf{B}$ (bimodularity embeddings). (B) Aggregation of embeddings at the level of graph nodes to build a feature vector that represents the sending and receiving partition of the source ($i$) and target ($j$) node respectively. (C) Feature matrix with features of all graph edges. (D) Edge feature space (first two SVD components). Each dot represents an edge of the graph colored based on its corresponding cluster. (E) Graph adjacency matrix where each edge cluster is colored. Sending (left) and receiving (top) nodes are shown with colors of the corresponding cluster. (F) Bicommunities of the block-cycle graph with colors corresponding to the edge cluster. Colors and markers of nodes indicate whether a node belongs to the sending (red, square), receiving (blue, diamond) or both (circle, purple) community.
@@
Bimodularity-based embedding of the graph nodes provides insightful representations of network organization. However, an additional step is needed to effectively detect pairs of sending $C_{k}^{\text{out}}$ and receiving communities $C_{k}^{\text{in}}$, termed bicommunities.
@@
Figure 3: Singular value decomposition (SVD) of the modularity matrix built from the C. elegans wiring network of chemical synapses. (A) Spectrum of the singular values of the SVD. Colors indicate the sign of the singular value associated to assortative (red, $\mu>0$) or dissortative (blue, $\mu<0$) community structures. (B) Projection of the graph nodes (neurons) onto the left ($\mathbf{u}_{1}$, horizontal axis) and right ($\mathbf{v}_{1}$, vertical axis) singular vectors of the bimodular component associated with the largest singular value ($\mu_{1}=10.99$). Colors and shape of nodes indicate the type of the neurons (sensory: red square; interneuron: blue circle; motor: green diamond; or sex specific: purple triangle). Edge colors indicate the type of the target node with the same color code as for the neuron types. Neuron labels are shown for graph nodes that are projected far from the origin $(0,0)$ of the axes.
@@
Figure 4 shows the 3 bicommunities of the C. elegans with highest bimodularity index (out of 5) extracted from the 5 first components of the SVD of $\mathbf{B}$. These clusters highlight specific information pathways in the worm’s motion system. In detail, $C_{3}$ identifies communication from body sensory processes to key interneurons in the worm’s brain (AVA, AVB, AVD). $C_{4}$ describes feed-forward transmissions of information from these interneurons to motor neurons of the body and interneurons of the tail (PVC). $C_{2}$ finally shows interconnectivity between motor neurons of the body and the neck as well as their communication with interneurons of the worm’s tail. Finer definitions of the C. elegans’ bicommunities is achieved with a larger number of edge clusters, see SI Appendix, Fig. S5.
Figure 4: Bicommunity of the C. elegans extracted from the $N=5$ first bimodularity embeddings (components of the SVD). Only the clusters with top 3 highest bimodularity index (out of $K=5$) are displayed. (A) The sending and receiving parts of each bicommunity are highlighted on the spatial location of neurons. Markers indicate whether a neuron belongs to the sending (red square), the receiving (blue diamond) or both (pink circle) parts of a bicommunity. The colors represent the difference between the proportion of edges in the sending and in the receiving partitions with red and blue for nodes that tend to send and receive more respectively. (B) The distribution of neuron types (sensory, inter, motor and sex-specific neurons) in each sending or receiving community is summarized in the pie charts (middle). (C) The spatial distribution of bicommunities along the tail to head axis is detailed in the histograms (right) as a proportion of the total number of neurons (gray bars). The sending and receiving patterns are separated in the upper and lower bars respectively. Colors indicate the type of neuron following the color scheme of the pie charts. The body and tail histograms are made bigger (twice their original size/number of neurons) for visualization purpose.
@@ Discussion
Methods previously tackled the embedding of directed graph nodes through the SVD of graph operators (e.g. adjacency or Laplacian) (29, 9, 28, 27). These approaches however usually fail to provide an interpretation of the embedding space. This limitation is further emphasized as they do not leverage the intrinsic correspondence between the left and right singular vectors through the associated singular value. The proposed bimodularity perspective is novel as it intuitively describes the relationship between a sending and its corresponding receiving projection. Then, the bimodular edge embedding is a natural approach to identify edges mapping from nodes with similar sending patterns to nodes with similar receiving ones. In comparison to conventional approaches, the proposed edge embedding is novel and unique to bimodularity in the interpretability of edge features (20, 21).
[…]
[id=AC] Advances in machine learning have recently also contributed to directed community detection (32, 33, 34). Specifically, graph neural networks that incorporate the adjacency line graph (i.e., edges of the original graph are turned into nodes and then connected when sharing an original node) (32), or a spectral convolution layer via the magnetic Laplacian that is Hermitian (33, 35), have proven to be promising avenues for learning node embeddings. However, learning-based methods rely on training processes in high-dimensional parameter spaces that limit their generalizability and interpretability. In addition, it is not straightforward to understand how directionality is exploited to represent communities in terms of sending or receiving constituents. To overcome this shortcoming, a shallower neural architecture has been proposed to learn node embeddings that represent the source and target roles (34). Specifically, single-layer neural models were used to extract embeddings from features of an alternating random walk. Such identification of source and target roles is similar to our concept of sending and receiving communities, albeit there is no correspondence between sending and receiving patterns, as is the case for bimodularity. These frameworks primarily focus on representation learning instead of unsupervised discovery. In contrast, SVD-based edge embedding and detection of bicommunities flow naturally from the new definition of bimodularity. Nevertheless, future work could integrate these operations into learning frameworks as well.
@@
The application of the proposed scheme to the C. elegans’ neuronal wiring diagram further confirmed its relevance. In particular, the embedding according to the first component of the SVD encompasses the organization of the worm’s head and body locomotion system and untangles information pathways such as head-to-head, body-to-body and head-to-body sensory-motor flows. Specifically, this latter communication scheme is mainly driven by the AVE neurons that could integrate information from the sensory/control neurons of the head system and further transmit signals to the body control and motor neurons. We also observed that most body motor neurons also send information to more anterior motor neurons in the neck and the head, which confirms previous findings of intricate feedforward loop between the head and body motion systems (36, 37, 38). In short, the innovative aspect of bimodularity is that key clusters of neurons and their projections are captured in a single bimodular component (i.e., by a left/sending and right/receiving singular vectors) of the directed modularity matrix.
@@
There is one methodological aspect that needs to be addressed when dealing with the SVD; i.e., the sign ambiguity singular values and vectors (39). Indeed, in the $\mathbf{B=U\Sigma V}^{T}$ decomposition the sign of a singular value and one singular vector: $\mathbf{B=U\Sigma V}^{T}=\mathbf{U(-\Sigma)(-V}^{T})$ can be freely swapped , which raises the need for a clear sign convention for univocal interpretation. In the case of undirected graph, the modularity matrix is symmetric and the singular value (eigenvalue) hints whether a community structure is assortative (positive) or rather dissortative (negative, also known as anti-modular) (14, 15, 40). Here, we propose a sign convention associated to the dot product between the two singular vectors (i.e., cosine of the angle between them) and show that it is related to the assortative (or dissortative) nature of the partitions, see SI Appendix, Fig. S7. In particular, we observe that the bimodularity index of a bicommunity is negative when there are more edges between the sending and receiving sets of nodes than within them. Therefore, this sign convention provides additional information about the observed community structure. In the C. elegans wiring network, the second component with the largest singular value is dissortative and supports the findings of the first component (Fig. 3) by highlighting head interneurons that densely send information to body motor neurons without necessarily being connected (see SI Appendix, Fig. S8 for detailed description).
@@
While bimodularity is a suitable metric for highlighting community structure in directed networks, as an extension of modularity it shares some of its limitations such as the resolution limit (41). Future work should investigate how this limit is influenced by directed information. In Supplementary Material (SI Appendix, Fig. S9), we present a number of variations on the generated graphs that confirm the relevance and robustness of the bimodularity and bicommunity detection. The presented method for bicommunity detection is parametrized by the number of bimodular components $N$ and the number of desired clusters $K$. While $K$ is here tied to the clustering algorithm (i.e., k-means), we show that influence of $N$ is attenuated by multiplying entries in the edge feature (singular vectors $\mathbf{u}_{k}$ or $\mathbf{v}_{k}$) by their corresponding singular value $\mu_{k}$ (see SI Appendix, Fig. S2). This is a common approach in spectral clustering to incorporate the spectrum of eigen-/singular values to weight the matrix of eigen-/singular vectors that is then fed to clustering (30, 31).
@@
All code and data used in this article are available as open-source code at https://github.com/MIPLabCH/Bimodularity.
@@
A directed binary graph was built from the wiring diagram of the C. elegans’ nervous system using information about 2194 chemical synapses between 279 neurons (graph nodes) (1). Out of all these connections, 1961 (89%) were asymmetric while only 233 (11%) were bidirectional. The undirected electric gap junction network was not used in the main analyses, but the bimodularity optimization and singular vector embedding of the joint (chemical synapses and gap junctions) graph are shown in SI Appendix, Fig. S10. Details about the type of each neuron (sensory, inter- or motor neurons) and their position in the worm’s body have been gathered from (38, 42) and are summarized in SI Dataset S1.
@@
$Q_{\text{bi}}$ $=\frac{1}{m}\sum_{k=1}^{K}\sum_{\begin{subarray}{c}i\in C_{k}^{\text{out}}\\ j\in C_{k}^{\text{in}}\end{subarray}}\left[A_{ij}-\mathbb{E}(A_{ij}|\mathcal{H}_{0})\right]$
@@
We consider the graph partition problem to separate the graph into two communities of nodes (sending community) that have common targets (receiving community). Let $\mathbf{s}^{\text{out}}$ be a separator vector that takes value $s_{i}^{\text{out}}=1$ if node $i$ is in one sending partition and $s_{i}^{\text{out}}=-1$ if it is in the other. Similarly, we define $\mathbf{s}^{\text{in}}$ as a separator vector for the receiving partition corresponding to the sending one. Therefore, $\delta_{C^{\text{out}},C^{\text{in}}}(i,j)$ can be rewritten as
$\frac{s_{i}^{\text{out}}s_{j}^{\text{in}}+1}{2}$
that is $1$ if $s_{i}^{\text{out}}=s_{j}^{\text{in}}$ and 00 otherwise. We then develop the expression of bimodularity as:
@@
$=\frac{1}{2m}\sum\limits_{i,j}\left(\left[A_{ij}-\frac{k_{i}^{\text{out}}k_{j}^{\text{in}}}{m}\right]s_{i}^{\text{out}}s_{j}^{\text{in}}+\left[A_{ij}-\frac{k_{i}^{\text{out}}k_{j}^{\text{in}}}{m}\right]\right).$
Knowing that $\sum\limits_{i,j}A_{ij}=m$, that $\sum\limits_{i}k_{i}^{\text{out}}=m$ and that $\sum\limits_{j}k_{j}^{\text{in}}=m$, the right part of the sum is equal to zero and the bimodularity index is can be expressed as a matrix multiplication:
@@
Maximization of bimodularity $Q_{\text{bi}}(\mathbf{s}^{\text{out}},\mathbf{s}^{\text{in}})$ is achieved by convex relaxation of the separator vectors $\mathbf{s}^{\text{out}},\mathbf{s}^{\text{in}}$. In essence, the separator vectors are not anymore restricted to +1/1+1/-1+ 1 / - 1 but under the constraint to be unitary vectors such that $(\mathbf{s}^{\text{out}})^{T}\mathbf{s}^{\text{out}}=1$ and $(\mathbf{s}^{\text{in}})^{T}\mathbf{s}^{\text{in}}=1$. We then write the unconstrainted optimization problem using Lagrange multipliers $\mu^{\text{out}}$ and $\mu^{\text{in}}$:
@@
Multiplying the first expression by $\mathbf{B}^{T}$, we then obtain $\mu^{\text{out}}\mathbf{B}^{T}\mathbf{s}^{\text{out}}=\mathbf{B}^{T}\mathbf{Bs}^{\text{in}}$, where we can substitute the second expression to obtain
$\mu^{\text{out}}\mu^{\text{in}}\mathbf{s}^{\text{in}}=\mathbf{B}^{T}\mathbf{Bs}^{\text{in}}.$
Here we recognize that $\mathbf{s}^{\text{in}}$ is an eigenvector of the symmetric matrix $\mathbf{B}^{T}\mathbf{B}$ with the corresponding eigenvalue $\lambda^{\text{in}}=\mu^{\text{out}}\mu^{\text{in}}$. Similarly, applying the same approach to the second (lower) expression and substituting the upper one, we obtain
$\mu^{\text{out}}\mu^{\text{in}}\mathbf{s}^{\text{out}}=\mathbf{BB}^{T}\mathbf{s}^{\text{out}},$
which shows that $\mathbf{s}^{\text{out}}$ is an eigenvector of $\mathbf{B}\mathbf{B}^{T}$ with the eigenvalue $\lambda^{\text{out}}=\mu^{\text{out}}\mu^{\text{in}}$.